OpenStudy (anonymous):

sqrt{(1-sqrt{x})/(1+sqrt{x})}

OpenStudy (anonymous):

\[\sqrt{(1-\sqrt{x})/(1+\sqrt{x})} \]

OpenStudy (anonymous):

\[\int\limits_{}^{}\sqrt{(1-\sqrt{x})/(1+\sqrt{x})}\]

OpenStudy (anonymous):

I would start by substituting u=sqrtx.

OpenStudy (anonymous):

Oh wait!! I didn't see the big square root =).

OpenStudy (anonymous):

Using the formula (a+b) (a-b) = a^2 - b^2 You get 1 - x under the big root sign

OpenStudy (anonymous):

Sorry, It has a divide sign !!!

OpenStudy (anonymous):

You can rationalise by multiplying the number under the big root sign by \[1 - \sqrt{x}\] -------------- \[1 - \sqrt{x}\]